# HPSS: Heuristic Prompting Strategy Search for LLM Evaluators

### v1

@@ Abstract
Since the adoption of large language models (LLMs) for text evaluation has become increasingly prevalent in the field of natural language processing (NLP), a series of existing works attempt to optimize the prompts for LLM evaluators to improve their alignment with human judgment. However, their efforts are limited to optimizing individual factors of evaluation prompts, such as evaluation criteria or output formats, neglecting the combinatorial impact of multiple factors, which leads to insufficient optimization of the evaluation pipeline. Nevertheless, identifying well-behaved prompting strategies for adjusting multiple factors requires extensive enumeration. To this end, we comprehensively integrate 8 key factors for evaluation prompts and propose a novel automatic prompting strategy optimization method called Heuristic Prompting Strategy Search (HPSS). Inspired by the genetic algorithm, HPSS conducts an iterative search to find well-behaved prompting strategies for LLM evaluators. A heuristic function is employed to guide the search process, enhancing the performance of our algorithm. Extensive experiments across four evaluation tasks demonstrate the effectiveness of HPSS, consistently outperforming both human-designed evaluation prompts and existing automatic prompt optimization methods.
@@ 1 Introduction
Evaluation is a long-standing critical and challenging task in NLP Celikyilmaz et al. (2020); Chang et al. (2024). Recently, with the advent of advanced large language models (LLMs) such as GPT-4, LLM-based evaluators have been widely used for various natural language generation tasks Wang et al. (2023a); Liu et al. (2023); Zheng et al. (2023); Hu et al. (2024). Leveraging the LLMs’ strong language understanding and instruction-following capabilities, the evaluation task can be addressed via an evaluation prompt that includes a task description, evaluation rules, and the text to be assessed Chen et al. (2023); Ke et al. (2023). Various studies have reported that LLMs can provide accurate and interpretable evaluation results of text quality on the specified aspect, offering stronger generalization ability compared with previous evaluation methods Papineni et al. (2002); Lin (2004); Zhang et al. (2020); Zhong et al. (2022).
@@
However, we argue that these methods are only limited to optimizing a single factor of evaluation prompts (e.g., evaluation criteria or output formats) but neglect the prompting strategy for adjusting multiple factors (Kim et al., 2023). Given that the evaluation prompts comprise multiple components Gao et al. (2024), each encompassing various factors that simultaneously affect the performance of LLM evaluators, it is necessary to collectively adjust these factors to fully harness its evaluation capability. Nevertheless, due to the vast state space involved in making these adjustments, extensive enumeration is required to identify well-behaved prompting strategies, highlighting the need for effective automatic optimization methods.
To this end, we comprehensively integrate a series of prompting strategies for LLM evaluators from previous works, select 8 key factors, and propose the Heuristic Prompting Strategy Search (HPSS) algorithm to automatically optimize prompting strategies for LLM evaluators tailored to specific evaluation aspects. Inspired by the genetic algorithm Mitchell (1980), HPSS gradually mutates the previous prompting strategies and selects candidate strategies based on their performance on a validation dataset. Leveraging the limited number of values within each factor, HPSS introduced a heuristic function to guide the search process, assigning higher exploration probability to more promising values, which enhances the effectiveness of mutation and leads to better performance compared to random search. We optimize the prompting strategy for two different LLM evaluators (i.e., GPT-4o-mini OpenAI (2023) and Qwen2.5-14B-Instruct Yang et al. (2024)) across four evaluation tasks for natural language generation (NLG). Experiment results show that HPSS achieves substantial improvement over the baseline of human-designed evaluation prompts and existing automatic prompt optimization methods. Compared to the commonly-used evaluation prompt from MT-Bench Zheng et al. (2023), HPSS can achieve an average 29.4% relative performance improvement with the same generation times. Furthermore, compared to other manually designed prompting strategies such as G-Eval Liu et al. (2023) and CloserLook Chiang and Lee (2023), HPSS can still achieve significantly better performance with only 5% of the generation times. Our contributions can be summarized as follows:
@@ 2.2 Prompt Optimization
Reference The reference answer for the task in the sample to evaluate. In addition to using no reference Liu et al. (2023), previous work also utilizes the following types of references: (1) Self-Generated Reference: prompting the evaluator to generate a reference independently and then feed it into $\mathcal{T}$ Zeng et al. (2024), (2) Dialectic: prompting the evaluator to generate a reference along with the final rating He et al. (2024) No Reference Self-Generated Reference Dialectic
@@
Although LLMs have demonstrated impressive performance across various NLP tasks Zhao et al. (2023), researchers point out that their abilities are highly dependent on sophisticated prompt design Pryzant et al. (2023); Leidinger et al. (2023); Raina et al. (2024). Considering that manual prompt engineering is time-consuming, many works attempt to optimize prompt automatically. GPS Xu et al. (2022) and GRIPS Prasad et al. (2023) employ a genetic algorithm, using rule or generative language models to mutate the initial prompt. PromptBreeder Fernando et al. (2023) and EvoPrompt Guo et al. (2023) utilize LLMs for mutation and crossover operations in evolutionary searches. OPRO Yang et al. (2023) demonstrates the potential of LLMs to perform iteratively prompt optimization based on search history. Furthermore, Hsieh et al. (2024) extends prompt optimization to long prompts containing multiple sentences. They utilized Lin-UCB to guide the selection of sentences and employed in-context learning for LLM-based mutation. For comparison, our work shifts the focus from word or sentence-level prompt optimization to a higher-level strategy that adjusts multiple factors of evaluation prompts. Additionally, we introduce a heuristic function to enhance the effectiveness of mutation based on the characteristics of evaluation prompting strategies.
@@ 3 Methodology
Figure 2: Overview of our HPSS algorithm. By perturbing the value of each factor in the prompting strategy from MT-Bench, HPSS gets the initial strategy population and the advantage estimations of values. Subsequently, new strategies are iteratively searched based on the guidance of the value advantage. The performance of new strategies is used to update the top-$k$ strategy population and the advantage estimation.
@@ 3.1 Problem Formulation
where $r_{D,a}(\mathcal{T})$ represents the performance of $\mathcal{T}$ for the aspect $a$ on the dataset $D$, which is typically measured by the correlation between the evaluation results of LLM evaluators using $\mathcal{T}$ and human judgments. To aid the search for the optimal prompting strategy, we will use a validation dataset with human judgments to evaluate the performance of $\mathcal{T}$. Without loss of generality, we simplify $r_{D,a}(\mathcal{T})$ as $r(\mathcal{T})$ to denote the performance of $\mathcal{T}$ for a specific aspect on the validation dataset.
@@ 3.2 Factors of Evaluation Prompt
Considering a series of prompting strategies from previous works, we select 8 key factors for the prompting strategy search, as described in Table 1. Detailed prompts are provided in Appendix A.
The first 7 factors belong to 3 main components of the evaluation prompt template $\mathcal{T}$: (1) Task Description (TD), which encompasses the definition of the evaluation task and output format (including Scoring Scale and Chain-of-Thought), (2) Evaluation Rules (ER), which encompass the standard and steps for evaluation (including Evaluation Criteria and AutoCoT), and (3) Input Content (IC), which encompasses the input sample and sample-specific auxiliary information (including In-Context Example, Reference, and Metrics). Previous works adopt different orders for these 3 components. For instance, MT-Bench Zheng et al. (2023) uses the order TD $\rightarrow$ ER $\rightarrow$ IC, while AutoCalibrate Liu et al. (2024c) uses TD $\rightarrow$ IC $\rightarrow$ ER. We consider the order of these 3 components as the last factor and all 6 possible orders as the selection range. The overall size of the search space is 12,960. We also conduct a preliminary experiment to explore the effect of these factors on Appendix B. The results show that these factors significantly influence the performance of LLM evaluators, yet some findings diverge from previous works.
@@ 3.3 Heuristic Prompting Strategy Search
Inspired by the Genetic Algorithms (GA) Mitchell (1980), we propose Heuristic Prompting Strategy Search (HPSS) to effectively search prompting strategies to help LLM evaluators achieve better alignment with human judgment. The framework is shown in Figure 2. HPSS maintains a population of $k$ top-performing prompting strategies and mutates each strategy in this population to create new strategies. Their performance is measured by the human correlation between LLM evaluators and human judgment on the validation dataset. Different from traditional GA that randomly selects values for mutation with equal probability, HPSS leverages the limited number of values within each factor and calculates the expected correlation metric advantage of each value over the random selection of the corresponding factor. The advantage of value $f_{ij}$ for factor $F_{i}$ denotes as $A_{ij}$. Then, this advantage score serves as a heuristic function to guide mutation. This mechanism assigns higher selection probabilities to more promising values, reducing the cost caused by the blindly random search in GA. HPSS mainly contains two steps:
@@
This step aims to get the initial strategy population and advantage estimation of values. Starting with a commonly-used reference-free pointwise grading prompting strategy from MT-Bench, we modify the selection value for each factor separately to construct multiple strategies. After that, we calculate the initial advantage estimation of each value based on their performance and select $k$ top-performing strategies for the subsequent iterative search step.
@@
In each iteration, HPSS mutates each prompting strategy $\mathcal{T}^{c}$ within the population for $g$ times to generate new strategies, and then update the top-$k$ strategy population based on the performance of these new strategies. To enhance the effectiveness of mutation, HPSS takes overall advantage of all factor values as the performance estimation of one strategy and allocates more search resources to the strategies with higher overall advantages. Specifically, we utilize two types of mutations: exploration and exploitation, which have the probability $\rho$ and $1-\rho$ to be chosen for each mutation, respectively.
@@
$B_{\mathcal{T}}=A_{ij}-A_{ic_{i}}+\lambda\sqrt{\ln(t)/{M_{ij}}}$ (2)
@@
After reaching the maximum search step, HPSS returns the best-performing strategy on the validation dataset as the final result. Detailed implementation of these two steps is presented in Appendix H.
@@ 4.1 Experimental Setup
Tasks and Datasets
We evaluate HPSS on four pointwise grading evaluation tasks: Summeval Fabbri et al. (2021) for text summarization, Topical-Chat Gopalakrishnan et al. (2019) for dialogue generation, SFHOT / SFRES Wen et al. (2015) for data-to-text generation, and HANNA Chhun et al. (2022) for story generation. The Spearman ($\rho$) correlation coefficient between human judgments and LLM evaluations is adopted as the performance metric. We also validate HPSS on pairwise comparison benchmarks MT-Bench Zheng et al. (2023), AUTO-J Li et al. (2024), and LLMBar Zeng et al. (2024). The results are provided in Appendix I. More details about benchmarks and metrics can be found in Appendix D and E. Following HD-Eval Liu et al. (2024d), for each dataset, a 50% proportion is held out for testing, while the rest is applied for validation.
Baselines
We compare HPSS with three types of baselines: (1) Non-LLM Evaluators: This category includes BLEU-4 Papineni et al. (2002), BertScore Zhang et al. (2020), and UniEval Zhong et al. (2022). (2) Human-Designed LLM Evaluators: The prompting strategy from MT-Bench Zheng et al. (2023) stands as the starting point of searching. G-Eval Liu et al. (2023) integrates AutoCoT to enhance the performance of LLM evaluators. Chiang and Lee (2023) explores various evaluation schemes. We use its best setting analyze-rate, denoted as CloserLook, and attempt to further improve its performance by providing human evaluation examples (CloserLook + ICL). (3) Automatic Prompt Optimization for LLM Evaluators: We implement 4 strong prompt optimization baselines, including APE Zhou et al. (2023b), OPRO Yang et al. (2023), Greedy Prasad et al. (2023); Zhou et al. (2023a) and Stepwise-Greedy. Greedy is a widely used iterative search algorithm, where random perturbations are applied to the current strategy to generate multiple candidates at each iteration, and the best-performing one is retained for the next iteration. Stepwise-Greedy sequentially optimizes each factor, choosing the best-performing value in each step. More details can be found in Appendix F.
@@
Models and Configurations
We choose the representative closed-source model GPT-4o-mini OpenAI (2023) and open-source model Qwen2.5-14B-Instruct Yang et al. (2024) as evaluation models. For G-Eval and CloserLook, we use their default settings (with 20 generation times and the decoding temperature as 1.0). For other LLM evaluators, greedy search is used for reproducibility. We limit the computational budget to 71 evaluations on the validation dataset (including 21 evaluations during initiation in HPSS) for all prompt search methods (21 for Stepwise-Greedy as an exception, as their evaluation count is constant). Results of prompt search on Qwen2.5-14B-Instruct are averaged over 3 seeds and the standard deviation is provided. While for GPT-4o-mini, we report the results of one seed due to the API cost. More details including the cost of HPSS are in Appendix G.

### v2

@@ Abstract
Since the adoption of large language models (LLMs) for text evaluation has become increasingly prevalent in the field of natural language processing (NLP), a series of existing works attempt to optimize the prompts for LLM evaluators to improve their alignment with human judgment. However, their efforts are limited to optimizing individual factors of evaluation prompts, such as evaluation criteria or output formats, neglecting the combinatorial impact of multiple factors, which leads to insufficient optimization of the evaluation pipeline. Nevertheless, identifying well-behaved prompting strategies for adjusting multiple factors requires extensive enumeration. To this end, we comprehensively integrate 8 key factors for evaluation prompts and propose a novel automatic prompting strategy optimization method called Heuristic Prompting Strategy Search (HPSS). Inspired by the genetic algorithm, HPSS conducts an iterative search to find well-behaved prompting strategies for LLM evaluators. A heuristic function is employed to guide the search process, enhancing the performance of our algorithm. Extensive experiments across four evaluation tasks demonstrate the effectiveness of HPSS, consistently outperforming both human-designed evaluation prompts and existing automatic prompt optimization methods. Our code is available at https://github.com/thu-coai/HPSS.
@@ 1 Introduction
Evaluation is a long-standing critical and challenging task in NLP Celikyilmaz et al. (2020); Chang et al. (2024). Recently, with the advent of advanced large language models (LLMs) such as GPT-4, LLM-based evaluators have been widely used for various natural language generation tasks Wang et al. (2023a); Liu et al. (2023); Zheng et al. (2023); Hu et al. (2024a); Ke et al. (2024). Leveraging the LLMs’ strong language understanding and instruction-following capabilities, the evaluation task can be addressed via an evaluation prompt that includes a task description, evaluation rules, and the text to be assessed Chen et al. (2023); Ke et al. (2023). Various studies have reported that LLMs can provide accurate and interpretable evaluation results of text quality on the specified aspect, offering stronger generalization ability compared with previous evaluation methods Papineni et al. (2002); Lin (2004); Zhang et al. (2020); Zhong et al. (2022).
@@
However, we argue that these methods are only limited to optimizing a single factor of evaluation prompts (e.g., evaluation criteria or output formats) but neglect the prompting strategy for adjusting multiple factors (Kim et al., 2023). Given that the evaluation prompts comprise multiple components Gao et al. (2024), each encompassing various factors that simultaneously affect the performance of LLM evaluators, it is necessary to collectively adjust these factors to fully harness its evaluation capability.
Nevertheless, due to the vast state space involved in making these adjustments, extensive enumeration is required to identify well-behaved prompting strategies, highlighting the need for effective automatic optimization methods.
To this end, we comprehensively integrate a series of prompting strategies for LLM evaluators from previous works, select 8 key factors, and propose the Heuristic Prompting Strategy Search (HPSS) algorithm to automatically optimize prompting strategies for LLM evaluators tailored to specific evaluation aspects. Inspired by the genetic algorithm Mitchell (1980), HPSS gradually mutates the previous prompting strategies and selects candidate strategies based on their performance on a validation dataset. Leveraging the limited number of values within each factor,
HPSS introduced a heuristic function to guide the search process, assigning higher exploration probability to more promising values, which enhances the effectiveness of mutation and leads to better performance compared to random search. We optimize the prompting strategy for two different LLM evaluators (i.e., GPT-4o-mini OpenAI (2023) and Qwen2.5-14B-Instruct Yang et al. (2024a)) across four evaluation tasks for natural language generation (NLG). Experiment results show that HPSS achieves substantial improvement over the baseline of human-designed evaluation prompts and existing automatic prompt optimization methods. Compared to the commonly-used evaluation prompt from MT-Bench Zheng et al. (2023), HPSS can achieve an average 29.4% relative performance improvement with the same generation times. Furthermore, compared to other manually designed prompting strategies such as G-Eval Liu et al. (2023) and CloserLook Chiang and Lee (2023), HPSS can still achieve significantly better performance with only 5% of the generation times. Our contributions can be summarized as follows:
@@ 2.2 Prompt Optimization
Reference The reference answer for the task in the sample to evaluate. In addition to using no reference Liu et al. (2023), previous work also utilizes the following types of references: (1) Human-Written Reference [^1] Constructing high-quality human-written references requires extensive manual annotation. We do not consider this usage in our study to ensure a fair comparison. Zheng et al. (2023); Doddapaneni et al. (2024), (2) Self-Generated Reference: prompting the evaluator to generate a reference independently and then feed it into $\mathcal{T}$ Zeng et al. (2024), (3) Dialectic: prompting the evaluator to generate a reference along with the final rating He et al. (2024) No Reference Self-Generated Reference Dialectic
@@
Although LLMs have demonstrated impressive performance across various NLP tasks Zhao et al. (2023), researchers point out that their abilities are highly dependent on sophisticated prompt design Pryzant et al. (2023); Leidinger et al. (2023); Raina et al. (2024). Considering that manual prompt engineering is time-consuming, many works attempt to optimize prompt automatically. For instance, GPS Xu et al. (2022) and GRIPS Prasad et al. (2023) employ a genetic algorithm, using rule or generative language models to mutate the initial prompt. PromptBreeder Fernando et al. (2024) and EvoPrompt Guo et al. (2024) utilize LLMs for mutation and crossover operations in evolutionary searches. OPRO Yang et al. (2024b) demonstrates the potential of LLMs to perform iteratively prompt optimization based on search history. Furthermore, Hsieh et al. (2024) extends prompt optimization to long prompts containing multiple sentences. They utilized Lin-UCB to guide the selection of sentences and employed in-context learning for LLM-based mutation. For comparison, our work shifts the focus from word or sentence-level prompt optimization to a higher-level strategy that adjusts multiple factors of evaluation prompts. Additionally, we introduce a heuristic function to enhance the effectiveness of mutation based on the characteristics of evaluation prompting strategies.
@@ 3 Methodology
Figure 2: Overview of our HPSS algorithm. By perturbing the value of each factor in the prompting strategy from MT-Bench, HPSS gets the initial strategy population and the advantage estimations of values. Subsequently, new strategies are iteratively searched based on the guidance of the value advantage. The performance $r$ of new strategies is typically measured by the correlation between the evaluation results of LLM evaluators and human judgments, which is used to update the top-$k$ strategy population and the advantage estimation.
@@ 3.1 Problem Formulation
where $r_{D,a}(\mathcal{T})$ represents the performance of $\mathcal{T}$ for the aspect $a$ on the dataset $D$, which is typically measured by the correlation between the evaluation results of LLM evaluators using $\mathcal{T}$ and human judgments. Detailed calculation of $r_{D,a}(\mathcal{T})$ can be found in Appendix A. To aid the search for the optimal prompting strategy, we will use a validation dataset with human judgments to evaluate the performance of $\mathcal{T}$. Without loss of generality, we simplify $r_{D,a}(\mathcal{T})$ as $r(\mathcal{T})$ to denote the performance of $\mathcal{T}$ for a specific aspect on the validation dataset.
@@ 3.2 Factors of Evaluation Prompt
Considering a series of prompting strategies from previous works, we select 8 key factors for the prompting strategy search, as described in Table 1. Detailed prompts are provided in Appendix B.
The first 7 factors belong to 3 main components of the evaluation prompt template $\mathcal{T}$: (1) Task Description (TD), which encompasses the definition of the evaluation task and output format (including Scoring Scale and Chain-of-Thought), (2) Evaluation Rules (ER), which encompass the standard and steps for evaluation (including Evaluation Criteria and AutoCoT), and (3) Input Content (IC), which encompasses the input sample and sample-specific auxiliary information (including In-Context Example, Reference, and Metrics). Previous works adopt different orders for these 3 components. For instance, MT-Bench Zheng et al. (2023) uses the order TD $\rightarrow$ ER $\rightarrow$ IC, while AutoCalibrate Liu et al. (2024c) uses TD $\rightarrow$ IC $\rightarrow$ ER. We consider the order of these 3 components as the last factor and all 6 possible orders as the selection range. The overall size of the search space is 12,960. We also conduct a preliminary experiment to explore the effect of these factors on Appendix C. The results show that these factors significantly influence the performance of LLM evaluators, yet some findings diverge from previous works.
@@ 3.3 Heuristic Prompting Strategy Search
Inspired by the Genetic Algorithms (GA) Mitchell (1980), we propose Heuristic Prompting Strategy Search (HPSS) to effectively search prompting strategies to help LLM evaluators achieve better alignment with human judgment. The framework is shown in Figure 2. HPSS maintains a population of $k$ top-performing prompting strategies and mutates each strategy in this population to create new strategies. Their performance is measured by the human correlation between LLM evaluators and human judgments on the validation dataset. Different from traditional GA that randomly selects values for mutation with equal probability, HPSS leverages the limited number of values within each factor and calculates the expected correlation metric advantage of each value over the random selection of the corresponding factor. The advantage of value $f_{ij}$ for factor $F_{i}$ denotes as $A_{ij}$, which can be formally defined as:
$A_{ij}=\mathop{\mathbb{E}}_{\{F_{1},F_{2},\ldots,F_{n}\}\backslash\{F_{i}\}}[r(\mathcal{T}_{F_{1},F_{2},\ldots,f_{ij},\ldots,F_{n}})$ (2)
$-\frac{1}{m_{i}}\sum_{k=1}^{m_{i}}r(\mathcal{T}_{F_{1},F_{2},\ldots,f_{ik},\ldots,F_{n}})]$
Then, this advantage score serves as a heuristic function to guide mutation. This mechanism assigns higher selection probabilities to more promising values, reducing the cost caused by the blindly random search in GA. HPSS consists of two steps:
@@
This step aims to get the initial strategy population and advantage estimation of values. Starting with a commonly-used reference-free pointwise grading prompting strategy from MT-Bench, we modify the selection value for each factor separately to construct multiple strategies. After that, we calculate the initial advantage estimation of each value based on its performance and select $k$ top-performing strategies for the subsequent iterative search step.
@@
In each iteration, HPSS mutates each prompting strategy $\mathcal{T}^{c}$ within the population for $g$ times to generate new strategies, and then updates the top-$k$ strategy population based on the performance of these new strategies. To enhance the effectiveness of mutation, HPSS takes overall advantage of all factor values as the performance estimation of one strategy and allocates more search resources to the strategies with higher overall advantages. Specifically, we utilize two types of mutations: exploration and exploitation, which have the probability $\rho$ and $1-\rho$ to be chosen for each mutation, respectively.
@@
$B_{\mathcal{T}}=\underbrace{A_{ij}-A_{ic_{i}}}_{\text{advantage disparity}}+\underbrace{\lambda\sqrt{\ln(t)/{M_{ij}}}}_{\text{additional exploration term}}$ (3)
@@
After reaching the maximum search step, HPSS returns the best-performing strategy on the validation dataset as the final result. Detailed implementation of these two steps is presented in Appendix F.1.
@@ 4.1 Experimental Setup
Tasks and Datasets.
We evaluate HPSS on four pointwise grading evaluation tasks: Summeval Fabbri et al. (2021) for text summarization, Topical-Chat Gopalakrishnan et al. (2019) for dialogue generation, SFHOT / SFRES Wen et al. (2015) for data-to-text generation, and HANNA Chhun et al. (2022) for story generation. The Spearman ($\rho$) correlation coefficient between human judgments and LLM evaluations is adopted as the performance metric. We also validate HPSS on pairwise comparison benchmarks MT-Bench Zheng et al. (2023), AUTO-J Li et al. (2024), and LLMBar Zeng et al. (2024). The results are provided in Appendix G. More details about benchmarks and metrics can be found in Appendix E and A. Following HD-Eval Liu et al. (2024d), for each dataset, a 50% proportion is held out for testing, while the rest is applied for validation.
Baselines.
We compare HPSS with three types of baselines: (1) Non-LLM Evaluators: This category includes BLEU-4 Papineni et al. (2002), BertScore Zhang et al. (2020), and UniEval Zhong et al. (2022). (2) Human-Designed LLM Evaluators: The prompting strategy from MT-Bench Zheng et al. (2023) stands as the starting point of searching. G-Eval Liu et al. (2023) integrates AutoCoT to enhance the performance of LLM evaluators. Chiang and Lee (2023) explores various evaluation schemes. We use its best setting analyze-rate, denoted as CloserLook, and attempt to further improve its performance by providing human evaluation examples (CloserLook + ICL). (3) Automatic Prompt Optimization for LLM Evaluators: We implement 4 strong prompt optimization baselines, including APE Zhou et al. (2023b), OPRO Yang et al. (2024b), Greedy Prasad et al. (2023); Zhou et al. (2023a) and Stepwise-Greedy. Greedy is a widely used iterative search algorithm, where random perturbations are applied to the current strategy to generate multiple candidates at each iteration, and the best-performing one is retained for the next iteration. Stepwise-Greedy sequentially optimizes each factor, choosing the best-performing value in each step. More details can be found in Appendix F.2.
@@
Models and Configurations.
We choose the representative closed-source model GPT-4o-mini OpenAI (2023) and open-source model Qwen2.5-14B-Instruct Yang et al. (2024a) as evaluation models. For G-Eval and CloserLook, we use their default settings (with 20 generation times and the decoding temperature as 1.0). For other LLM evaluators, greedy search is used for reproducibility. We limit the computational budget to 71 evaluations on the validation dataset (including 21 evaluations during initiation in HPSS) for all prompt search methods (21 for Stepwise-Greedy as an exception, as their evaluation count is constant). Results of prompt search on Qwen2.5-14B-Instruct are averaged over 3 seeds and the standard deviation is provided. While for GPT-4o-mini, we report the results of one seed due to the API cost. More details including the cost of HPSS are in Appendix F.3.